OpenStudy (anonymous):

In 0.0775 Find the natural logarithm, base e.

OpenStudy (anonymous):

i don't understand... you want the approximate value? use a calculator...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you have a calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how about the calculator on your computer?

OpenStudy (anonymous):

I'm not good with logarithm.

OpenStudy (anonymous):

suppose it is x, x^e=.0775

OpenStudy (anonymous):

now use calculator as said by dpalnc

OpenStudy (campbell_st):

use change of base... if you have base e log and want a base 10 log \[\log _{10}=\frac{\log _{10}(0.0775)}{\log _{10}e}\]

OpenStudy (anonymous):

-2.5575

OpenStudy (campbell_st):

I've done the calculation... and agree with your answer